OpenStudy (anonymous):

prove that this is an identity... sin^2x(cotx+1)^2=cos^2x(tanx+1)^2 Please help, these are really confusing.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

I have it too

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

@hero does it usually take this long to get the solution ? I got a bunch of these to do =/

hero (hero):

There's a specific approach to this that works most simply. First apply \(a^2b^2 = (ab)^2\) \(\sin^2x(\cot(x) + 1)^2 = [\sin(x)(\cot(x) + 1]^2\)

hero (hero):

No, it shouldn't take too long if you are very familiar with these. I already explained to you that my instructor allowed me to do these a different way so I had a different experience than most students.

hero (hero):

Okay so now we have \[[\sin(x)(\cot(x) + 1)]^2\] Change \(\cot(x)\) to \(\dfrac{\cos(x)}{\sin(x)}\)

hero (hero):

\[\left[\sin(x)\left(\frac{\cos(x)}{\sin(x)} + 1\right)\right]^2\] Factor out the \(\sin(x)\) in the denominator to get \[\left[\frac{\sin(x)}{\sin(x)}\left(\cos(x) + \sin(x)\right)\right]^2\] so now you have just \[[\cos(x) + \sin(x)]^2\]

hero (hero):

But now factor out \(\cos(x)\) to get: \[\left[\cos(x)\left(1 + \frac{\sin(x)}{\cos(x)}\right)\right]^2\]

hero (hero):

Change \(\dfrac{\sin(x)}{\cos(x)}\) to \(\tan(x)\) \[\left[\cos(x)\left(1 + \tan(x)\right)\right]^2\]

hero (hero):

And then re-distribute the squared to get: \[\cos^2(x)(1 + \tan(x))^2\] which is the same as \[\cos^2(x)(\tan(x) + 1)^2\]

OpenStudy (anonymous):

Thanks so much!